OpenStudy (anonymous):

when it says "an angle measure of the hexagon" which each individual angle is 120 degrees, does it want the sum of angle measures which is 720 degrees or just the individual?

ganeshie8 (ganeshie8):

does it say "an angle measure of the REGULAR hexagon" ? then yes it wants 120 degrees

ganeshie8 (ganeshie8):

default we take it as interior angle if it doesnt say explicitly its exterior angle

ganeshie8 (ganeshie8):

maybe if you can attach the screenshot of what u see there.. we can see what it says exactly..

OpenStudy (anonymous):

EXACTLY

ganeshie8 (ganeshie8):

yes the question is fully ambiguous

OpenStudy (anonymous):

grrr, atleast someone sees it, my teach is soo critical, its either 120 or 720 ill put both?

ganeshie8 (ganeshie8):

it doesnt say whether its regular hexagon, so we cant assume each angle is 120

ganeshie8 (ganeshie8):

i think ur teach wants u to find the x

OpenStudy (anonymous):

i could also put that x=17. very true.

ganeshie8 (ganeshie8):

if hmm yes.. .if we assume its regular, then both angles show wud be equal.. put something and discuss wid ur teacher... :)

ganeshie8 (ganeshie8):

or you can put (5x+35) or (8x-16) directly they're also angles from the way its shown in pix

OpenStudy (anonymous):

yesyessyes. thank you ganeshi!

ganeshie8 (ganeshie8):

np :) and wait, Directrix has somthing to add it seems

Directrix (directrix):

My best guess is that the question is intended to be the following: In a regular hexagon, the interior angles are congruent. If two angles of the hexagon have measures as shown, what is the measure of an interior angle of a regular hexagon. Note: It is not assumed that the student already knows that the angles of a regular hexagon have measure 120. 8x-15=5x+35 x = 17 8*17 -15 = 120 5*17 + 35 = 120 Answer: 120 One of the difficulties in answering an Asker's question is that the Answerer does not know the level of sophistication of study of the Asker.

ganeshie8 (ganeshie8):

haha yes ! i feel your teacher is super smart @dfgfdffdhfh . if it was a direct question, you wouldnt be exploring this much on hexagon... :D

OpenStudy (anonymous):

thanks so much @Directrix and @ganeshie8

Directrix (directrix):

We are happy to help.